OpenStudy (anonymous):

angle between v=(-7,8) w=(1,0)

OpenStudy (anonymous):

i don't get how you find v1 and v2

sam (.sam.):

Are those vectors

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i use v1*v2 right? but how do i get v1 and v2

sam (.sam.):

Use \[A • B=|A||B|\cos \theta\]

OpenStudy (anonymous):

what are A and B the sum of the vectors? the product of the vectors?

sam (.sam.):

\[\cos \theta=\frac{A•B}{|A||B|} \]